# The Dichotomous Nature of the $\sigma$ Meson and the Nucleon D-Term

### v1

@@ Abstract
Employing a symmetry-preserving contact-interaction formulation of the Dyson-Schwinger equations in quantum chromodynamics (QCD), we examine the identity of the $\sigma$ meson and its implications for the gravitational structure of hadrons. In this framework, the scalar meson emerges as the chiral partner of the pion, with both states’ properties tightly connected to the mechanisms of mass generation in QCD. We find that, above a critical coupling that triggers dynamical chiral symmetry breaking, the D-terms of the constituent quark, the pion, and $\sigma$ saturate at fixed values $D_{q,\pi,\sigma}=-1/3,-1,-7/3$. By examining the coupling strength evolution of the D-terms, this pattern follows naturally once a dual nature for the $\sigma$ meson is recognized: it behaves both as a quark–antiquark composite and as a dilaton arising from spontaneous scale symmetry breaking. This unified picture yields the prediction $D_{N}\sim-3$ for the nucleon D-term, consistent with contemporary lattice-QCD, continuum, and dispersive studies.
@@
In conventional quark models that account for dynamical CSB, such as the Nambu–Jona–Lasinio framework, [5, 4], the scalar–isoscalar channel arises naturally as the chiral partner of the pion, interpreted as a correlated exchange of pairs in the scattering process by Nambu. Nevertheless, in the real world, its exceptionally large width, [17, 10], and its prominent role in $\pi\pi$ scattering, e.g. [18, 19, 20], challenge a simple quark–antiquark interpretation. Furthermore, like the pion, the $\sigma$ is unusually light, with $m_{\sigma}\sim 400-500$ MeV. These features raise the question of whether this state might be a (pseudo) NG boson associated with a broken symmetry, [21, 22, 23, 24, 25]. The answer may lie in chiral effective theories, where the $\sigma$ emerges as the lightest scalar–isoscalar excitation and naturally connects to the trace anomaly and the gravitational form factors (GFF), e.g. [26, 27, 28]. The latter are actual probes of the energy-momentum tensor and thus scale symmetry [29, 30, 31, 32]. Closely related near-conformal scenarios similarly predict a light dilatonic scalar whose mass is controlled by the smallness of the QCD $\beta$-function in the infrared [33, 34, 35]. Moreover, recent lattice QCD explorations of the pion and nucleon GFFs, [36, 37], are found compatible with scalar meson dominance and a dilatonic view of the $\sigma$ meson [38, 28]. Altogether, these developments support a description of the $\sigma$ as the (pseudo)–NG boson associated with spontaneous scale symmetry breaking (SSB).
@@
Here $d$ is the spacetime dimension, $d_{\varphi}=(d-2)/2$ is the dimension of a free scalar field $\varphi$, and $c_{3}$ denotes coefficient of the cubic term in the dilaton potential. The quantity $\Delta_{q\bar{q}}$ represents the scaling dimension of the operator $q\bar{q}$, determined as $\Delta_{q\bar{q}}=d-2$ to ensure the soft pion theorem, [27]. Expressed in this way, the 1/(d1)-1/(d-1) and 0 terms on the the right hand side coincide precisely with the intrinsic quark contribution to the meson and quark D-terms at the critical coupling, while the remaining components arise from the dilaton pole contributions. One may even speculate that, in analogy with the quark case, that the nucleon D-term likewise receives no intrinsic contribution and is instead saturated by pole terms alone. This would align with the view that the fermion D-term is of purely dynamical origin [32, 31, 64].
@@
Notably, this seemingly naive estimate aligns with contemporary results from lattice QCD [37], continuum approaches [52], and dispersive analyses [65]. In particular, lattice QCD yields $D_{N}=-3.87(97)$ and 3.35(58)-3.35(58) from dipole and z-expansion fits, respectively; continuum studies report $D_{N}=-3.114(10)$ and dispersive methods find $D_{N}=-3.38_{-0.35}^{+0.34}$.
@@ I Acknowledgments
Work supported by: National Natural Science Foundation of China, grant no. 12135007; Postdoctoral Fellowship Program of CPSF under Grant Number GZC20240759; Spanish Ministry of Science and Innovation (MICINN grant no. PID2022-140440NB-C22); and Junta de Andalucía (grant no. P18-FR-5057).

### v2

@@ Abstract
Employing a symmetry-preserving contact-interaction formulation of the Dyson-Schwinger equations in quantum chromodynamics (QCD), we examine the identity of the $\sigma$ meson and its implications for the gravitational structure of hadrons. In this framework, the scalar meson emerges as the chiral partner of the pion, with both states’ properties tightly connected to the mechanisms of mass generation in QCD. We find that, above a critical coupling that triggers dynamical chiral symmetry breaking, the D-terms of the constituent quark, the pion, and $\sigma$ saturate at fixed values $D_{q,\pi,\sigma}=-1/3,-1,-7/3$. By examining the coupling strength evolution of the D-terms, this pattern follows naturally once a dual nature for the $\sigma$ meson is recognized: it behaves both as a quark–antiquark composite and as a dilaton arising from spontaneous scale symmetry breaking. This unified picture yields the prediction $D_{N}\sim-3$ for the nucleon D-term, consistent with lattice QCD, dispersive analyses, and diverse continuum studies.
@@
In conventional quark models that account for dynamical CSB, such as the Nambu–Jona–Lasinio framework, [5, 4], the scalar–isoscalar channel arises naturally as the chiral partner of the pion, interpreted as a correlated exchange of pairs in the scattering process by Nambu. Nevertheless, in the real world, its exceptionally large width, [17, 10], and its prominent role in $\pi\pi$ scattering, e.g. [18, 19, 20], challenge a simple quark–antiquark interpretation. Furthermore, like the pion, the $\sigma$ is unusually light, with $m_{\sigma}\sim 400-500$ MeV. These features raise the question of whether this state might be a (pseudo) NG boson associated with a broken symmetry, [21, 22, 23, 24, 25]. The answer may lie in chiral effective theories, [26], where the $\sigma$ emerges as the lightest scalar–isoscalar excitation and naturally connects to the trace anomaly and the gravitational form factors (GFF), e.g. [27, 28]. The latter are actual probes of the energy-momentum tensor and thus scale symmetry [29, 30, 31, 32]. Closely related near-conformal scenarios similarly predict a light dilatonic scalar whose mass is controlled by the smallness of the QCD $\beta$-function in the infrared [33, 34, 35]. Moreover, recent lattice QCD explorations of the pion and nucleon GFFs, [36, 37], are found compatible with scalar meson dominance, [38, 39, 28], and a dilatonic view of the $\sigma$ meson [28]. Altogether, these developments support a description of the $\sigma$ as the (pseudo)–NG boson associated with spontaneous scale symmetry breaking (SSB).
@@
Here $d$ is the spacetime dimension, $d_{\varphi}=(d-2)/2$ is the dimension of a free scalar field $\varphi$, and $c_{3}$ denotes coefficient of the cubic term in the dilaton potential. The quantity $\Delta_{q\bar{q}}$ represents the scaling dimension of the operator $q\bar{q}$, determined as $\Delta_{q\bar{q}}=d-2$ to ensure the soft pion theorem, [27, 34, 35]. Expressed in this way, the 1/(d1)-1/(d-1) and 0 terms on the the right hand side coincide precisely with the intrinsic quark contribution to the meson and quark D-terms at the critical coupling, while the remaining components arise from the dilaton pole contributions. One may even speculate that, in analogy with the quark case, that the nucleon D-term likewise receives no intrinsic contribution and is instead saturated by pole terms alone. This would align with the view that the fermion D-term is of purely dynamical origin [32, 31, 65].
@@
Notably, this seemingly naive estimate aligns with contemporary results from lattice QCD [37], continuum approaches [53], and dispersive analyses [66]. In particular, lattice QCD yields $D_{N}=-3.87(97)$ and 3.35(58)-3.35(58) from dipole and z-expansion fits, respectively; continuum studies report $D_{N}=-3.114(10)$ and dispersive methods find $D_{N}=-3.38_{-0.35}^{+0.34}$. Notably, the scalar meson dominance yields D=3.0(4)D=-3.0(4)[39], and the dilaton picture examined thoroughly in Ref. [28] supports a $\sigma$-meson contribution to the nucleon’s D-term of $D_{N}^{\sigma}=-3.01(39)$.
@@ I Acknowledgments
We thank Fei Gao and Roman Zwicky for valuable constructive comments on the manuscript. Work supported by: National Natural Science Foundation of China, grant no. 12135007; Postdoctoral Fellowship Program of CPSF under Grant Number GZC20240759; Spanish Ministry of Science and Innovation (MICINN grant no. PID2022-140440NB-C22); and Junta de Andalucía (grant no. P18-FR-5057).